# You Sound a Little Tense: L2 Tailored Clear TTS Using Durational Vowel Properties

### v1

@@ Abstract
We present the first text-to-speech (TTS) system tailored to second language (L2) speakers. We use duration differences between American English tense (longer) and lax (shorter) vowels to create a “clarity mode” for Matcha-TTS. Our perception studies showed that French-L1, English-L2 listeners the participants had fewer (at least 9.15%) transcription errors when using our clarity mode, and found it more encouraging and respectful than overall slowed down speech. Remarkably, listeners were not aware of these effects: despite the decreased word error rate in clarity mode, listeners still believed that slowing all target words was the most intelligible, suggesting that actual intelligibility does not correlate with perceived intelligibility. Additionally, we found that Whisper-ASR did not use the same cues as L2 speakers to differentiate difficult vowels and is not sufficient to assess the intelligibility of TTS systems for these individuals.
@@ 1 Introduction
[…]
@@ 3 L2 clarity TTS
Given the improved perception of tense/lax vowels with simple, linguistically-driven duration changes, we added this as a new ‘clarity mode’ in Matcha-TTS [24] for L2 speech.
To enable L2 clarity mode in Matcha-TTS, we added a clarity flag that can be set to “True” or “False” at synthesis, along with a markup to control which words are emphasized. The user (or large language model driving a dialogue system) surrounds difficult words with exclamation points, e.g., “!peel!”, allowing the TTS to parse the words to be treated for clarity through several steps: 1) Parse each flagged word to see if it contains a tense or lax vowel, 2) If the word contains tense vowels but no lax vowels, the clarity modification is applied to the tense vowel containing word, 3) If the word contains both tense and lax vowels, and if the tense vowel has primary stress, the clarity modification is applied to the tense vowel containing word.
The modifications are made by applying an array ($carray$) with the same length as the phonemized phrase to scale the predicted duration of each phoneme ($w$), a method shown to result in natural and effective duration changes in TTS systems containing a phonemizer [25]. This clarity duration multiplier is applied after the base speech rate multiplier, which is an array containing the speech rate provided at synthesis (in our case, 0.75 which in our pilot was found to have a natural, conversational speech rate for English L1 listeners). The predicted durations ($w$) result from the Hadamard product of the text encoder outputs (Equation 1): $\log w$: the log duration predicted by the duration predictor as in [26], and $x_{\text{mask}}$: the mask for the text input indicating which values are valid [27]. In Matcha-TTS this is then scaled to the input duration the $speechrate$ array. The resulting $y_{\text{lengths}}$ and $y_{\text{max\_length}}$ are then used to calculate the attention alignment map for the text encoder.
@@
[…]
@@ 3.1 Stimulus generation
We tested 4 different TTS styles: 1. Base: the base Matcha-TTS (0.75x speech rate), 2. Stretch: 1.2x (0.75×1.6)0.75$\times$ 1.6)0.75 × 1.6 ) speech rate applied across the entire phrase, 3. Emphasis: 1.6x stretch applied across all target words (0.75x speech rate elsewhere), and 4. Clarity: 1.6x stretch applied across the target words containing a tense vowel (0.75x speech rate elsewhere).
[…]
@@
Because participants would hear each of the phrases 4 times (each TTS style had the same phrases), we added 1 confusion phrase (randomly assigning a TTS style) for each of the phrases. A confusion phrase was the same phrase context using the opposite minimal pair, for example: “She kept mentioning cot during the conversation.” and the confusion phrase “She kept mentioning cut during the conversation.” Lastly, for the single target word TTS, “clarity” TTS was the same as “emphasis” for a single tense vowel target and “base” for a single lax vowel target. As such, we assessed the TTS in terms of the length of the target words rather than TTS styles and ensured no repeated phrases with the same treatment.
@@
iMOS 5.12 $<$.001 Lax: targetS/base,fullS p=.043, p=.008; Tense: targetS/base p=.016
nMOS 110.83 $<$.001 Lax: all/fullS p$<$.001 targetS/base p=.006; Tense: all/fullS p$<$.001
@@
pMOS 30.24 $<$.001 Lax: all/fullS p$<$.001 targetS/base p=.028; Tense: all/fullS p $<$.001
Enc. 21.45 $<$.001 Lax: all/fullS p$<$.001 targetS/base p=.036; Tense: all/Stretch p$<$.001
Resp. 17.90 $<$.001 Lax: targetS/fullS,base p$<$.001, p=0.022 base/fullS p=0.001; targetS/fullS p$<$.001 base/fullS p=0.001
@@ 3.2 Experimental procedure
We conducted an experiment to assess the objective (through word error rate) and subjective (through mean opinion scores) performance of French L1 English L2 listeners using our “clarity mode” compared to the baseline models (i.e., “base”, “stretch”, “emphasis”). In an online experiment, participants chose to receive instructions in either English or French. They then provided demographic information on the language they first learned, their most commonly used daily language, their age and gender, and their self-rated English proficiency. Each participant was randomly assigned to start in either the single- or double-word trial.
The participants were shown the phrase with the target words removed, e.g., “Write down the word X followed by the word X on the paper,” and could play the audio only once. They then selected which words they heard and were told (in the double-word case) that it was possible to hear the same word twice. It was never the case that they would hear the same word, however, this instruction was added to encourage participants not to base their choice for the second word on what they believe they heard in the first word. The missing word was selected from a list of 4 words: for the single-word trial, 2 words were the tense/lax vowel minimal pair (e.g., beat, bit); the third word was a minimal pair word with another vowel (e.g., bat); and the final word was dissimilar from the other three choices (e.g., shop). The dissimilar choice functioned as an attention check. In the double word case, when the two words were a minimal pair, the selection for the remaining two words was as per the single-word trial; if the two words were not a minimal pair, the 4 choices were the two words and their minimal pairs. The order of the phrases was randomized for each participant. The experiment set up can be seen in Fig. 1.
Once the participants selected which word they heard, they responded to a questionnaire containing the MOS-X2 [29] naturalness (nMOS), intelligibility (iMOS), and prosody (pMOS) scores, as well as an intelligibility question for listening effort (eMOS) (“Please rate the degree of effort you had to make to understand the message”) from MOS-X [29]. The listening effort question was added to understand if there is a difference between being able to understand the phrase and how much effort was required to understand the phrase. They then responded to two questions from L2-directed speech research [12]: “For an English second language speaker being spoken to with this voice, how respectful is the voice?” (Resp.: condescending - respectful), “For an English second language speaker being spoken to with this voice, how encouraging is the voice?” (Enc: not encouraging - encouraging). These questions were to help us understand if slowing down or adding emphasis makes the TTS sound condescending, as can be the case with L2-directed speech [11, 12]. All scores were on a 10-point Likert scale.
@@ 3.2.1 Participants
We recruited N=56 participants (29F, age = 34.59$\pm$10.67) via Prolific. French was our target L1 population and the participant language demographics were as follows: daily language = French (40), English (14), French and English (2), Italian (1); English proficiency (1-5) = 5 (29), 4 (20), 3 (5), 2 (2); 71.4% of participants chose to have the instructions in French. The study received internal ethics approval.
@@ 3.3.1 Single word
To confirm our previous findings, we expect improved performance in WER with lengthened tense vowels but decreased performance for the same treatment in lax vowels. We computed one-way ANOVAs (Type II) followed by a post-hoc Tukey HSD as well as word error rates (WER) as the proportion of incorrect target words for each type of target word treatment. In the single word case “Clarity” is the same as “Base” for lax vowel target words, and “Emphasis” for tense vowel target words. As such, we present the results in the three defined states existing for the single word case: “base” has the baseline duration on the target, “target stretch” is 1.6x the duration of the baseline on the target, and “full stretch” is 1.2x the duration on the entire phrase.
We observed that, through our “clarity” mode stretch applied to tense-vowel-containing words, we could overcome the bias towards lax vowels, i.e., the fact that participants were more likely to respond, for example, “pill” than “peel,” which we observed in our validation study. By lengthening the target tense-vowel-containing word, the WER on the baseline (“base tense”) was reduced from 60.23% to 29.48% for “target stretch tense” (Table 2, Fig. 2). Importantly, we also confirmed that, indeed, stretching lax vowels, as is typical in L2-directed speech, results in more errors (28.90% vs. 16.86%). Lastly, the “stretch” TTS resulted in a slight reduction in WER over the baseline for tense vowels, yet not as much as for only emphasizing the target word. This suggests that a difference between the target word and its context was important for determining vowel perception [19].
Surprisingly, despite the objective improvements seen in the WER, the participants perceived a “target stretch” significantly ($\alpha=0.05$) more intelligible for lax vowels (Tables 2 and 3, Figs. 3, 4). Moreover, “target stretch” also required less listening effort (although not significantly). Furthermore, although the “full stretch” TTS shows improvements in WER over the baseline for tense vowels, it was found to be significantly less respectful and encouraging in all cases. We also saw that the “full stretch” TTS was considered to be less natural and had worse prosody than the other TTS styles. This confirms our expectation that maintaining a natural (for L1 speakers) speech rate outside of difficult words can make the L2 listeners feel less patronized by the voice. Interestingly, the “target stretch” lax was rated as more natural and as having better prosody than “base” lax vowel, once again, despite have lower objective comprehension scores. These findings are particularly interesting as they suggest that L2 speakers may be overconfident in their ability to use the formant cues when the vowel is lengthened.
@@ 3.3.2 Double word
We observed the lowest WER with our proposed “clarity” TTS, both overall and for lax vowels, while vastly improving the performance for tense vowels over the baseline (relative 50% improvement) (Table 4, Fig. 5). Again, the “base” TTS showed a bias towards lax vowels (tWER: 30.00%, lWER: 18.66%), that we were able to overcome with our “clarity” mode. Additionally, we see that simply emphasizing all difficult words that had a tense/lax minimal pair decreases performance in lax vowels. Lastly, we observed “stretch” having improved performance over both “emphasis” and “base” for tense vowels. Yet, as in the “emphasis” condition, we saw participants struggle to understand the short, lax vowels. Although the tense vowel words between “emphasis” and “clarity” and lax vowel words between “clarity” and base had the same duration, “clarity” mode still had a lower WER when specifically comparing these words. This likely resulted from the phrases containing both a tense and a lax vowel, where the L2 participants could use duration differences between the two words to more easily differentiate the words. A more in-depth exploration of these differences remains a topic for future work.
We once again observed the fascinating result that the participants rate the “emphasis” TTS the highest in all categories (although in this case, the scores are not significantly higher than those for the “clarity” TTS) (Tables 5 and 6, and Fig. 6), despite the WER being higher for this TTS than both “clarity” and “stretch” TTS styles. We observed that “stretch” is less natural and has poorer prosody than all other TTS styles, despite showing objective improvements in WER over “emphasis” and “base”. Lastly, we found that L2 participants rated both speaking too fast (“base”) and too slow (“stretch”) as being less respectful and less encouraging.
@@
iMOS 20.15 $<$.001 all/Base p$<$.001
nMOS 20.15 $<$.001 all/Stretch p$<$.001; Emphasis,Clarity/Base p$<$.001
eMOS 24.15 $<$.001 all/Base p$<$.001
pMOS 77.15 $<$.001 all/Stretch $<$.001; Emphasis,Clarity/Base p$<$.001
Enc. 35.74 $<$.001 Emphasis,Clarity/Stretch,Base p$<$.001
Resp. 35.87 $<$.001 Emphasis,Clarity/Stretch, Base p$<$.001
@@ 3.3.3 Whisper ASR
Similar to L2 speakers, we saw for the “base” TTS, Whisper struggled to predict the correct target words that lack context in the phrase (21.4% vs an expected 5-10% WER for this model), although the performance was slightly higher than that for the L2 speakers (Table 7, Fig 7). We did not see the same improvements in WER with the “clarity” TTS that we saw in the L2 participants. Instead, we saw an overall slowing down (“stretch”) of the TTS decreases the WER in the target words. Yet, we also saw that while the WER on the target words decreased with this slowing down, the proportion of errors in the target word stemming from the minimal pair substitutions was much higher for the “base” TTS (71.42%, all other TTS $<$ 37%), while the overall difference in WER both on the whole phrase and the target words was within 3% for all TTS styles. This suggests that while slowing down slightly reduces the overall number of errors, the target words were being predicted as even further from the target, e.g. where “peel” was replaced with “pill” in the “base” TTS was replaced with “peaked” in the “stretch” TTS. Therefore, the ASR does not use the same duration mechanisms as humans when facing difficult predicting words.

### v2

@@ Abstract
We present the first text-to-speech (TTS) system tailored to second language (L2) speakers. We use duration differences between American English tense (longer) and lax (shorter) vowels to create a ``clarity mode'' for Matcha-TTS. Our perception studies showed that French-L1, English-L2 listeners the participants had fewer (at least 9.15%) transcription errors when using our clarity mode, and found it more encouraging and respectful than overall slowed down speech. Remarkably, listeners were not aware of these effects: despite the decreased word error rate in clarity mode, listeners still believed that slowing all target words was the most intelligible, suggesting that actual intelligibility does not correlate with perceived intelligibility. Additionally, we found that Whisper-ASR did not use the same cues as L2 speakers to differentiate difficult vowels and is not sufficient to assess the intelligibility of TTS systems for these individuals.
@@ 1 Introduction
To improve TTS for second language speakers, we present a clarity mode for English TTS, built upon a perception-based approach. L2 speakers of English, particularly those whose L1 does not have vowel tensity contrast, often struggle to differentiate English lax (less extreme articulation and shorter) vowels with tense (more extreme articulation and longer) vowels [13, 14, 15, 16, 17, 18]. For example, French speakers may replace the lax /I/ (ship) by the tense /i/ (sheep) [16]. In our paper [19], psychophysical reverse-correlation [20] was used to reconstruct L1-English and L2-English listener’s mental representation of duration and pitch in tense-lax vowel contrasts in a data-driven manner. For L2-English speakers who spoke French, Mandarin and Japanese as an L1, it was found that vowel duration, but not pitch, affected perception of tense/lax vowels. This suggests that we can apply a duration mechanism to improve comprehension of this English vowel contrast when an L2 listener struggles to use the primary formant cues [21]. Following these results, we create, for the first time, an L2 ``clarity" mode within a TTS to improve L2 comprehension of these difficult American English vowels.
@@ 3 L2 clarity TTS
Given the improved perception of tense/lax vowels with simple, linguistically-driven duration changes, we added this as a new `clarity mode' in Matcha-TTS [24] for L2 speech.
To enable L2 clarity mode in Matcha-TTS, we added a clarity flag that can be set to ``True'' or ``False'' at synthesis, along with a markup to control which words are emphasized. The user (or large language model driving a dialogue system) surrounds difficult words with exclamation points, e.g., ``!peel!'', allowing the TTS to parse the words to be treated for clarity through several steps: 1) Parse each flagged word to see if it contains a tense or lax vowel, 2) If the word contains tense vowels but no lax vowels, the clarity modification is applied to the tense vowel containing word, 3) If the word contains both tense and lax vowels, and if the tense vowel has primary stress, the clarity modification is applied to the tense vowel containing word.
The modifications are made by applying an array (carraycarray) with the same length as the phonemized phrase to scale the predicted duration of each phoneme ($w$), a method shown to result in natural and effective duration changes in TTS systems containing a phonemizer [25]. This clarity duration multiplier is applied after the base speech rate multiplier, which is an array containing the speech rate provided at synthesis (in our case, 0.75 which in our pilot was found to have a natural, conversational speech rate for English L1 listeners). The predicted durations ($w$) result from the Hadamard product of the text encoder outputs (Equation 1): $\log w$: the log duration predicted by the duration predictor as in [26], and $x_{\text{mask}}$: the mask for the text input indicating which values are valid [27]. In Matcha-TTS this is then scaled to the input duration the speechratespeechrate array. The resulting $y_{\text{lengths}}$ and $y_{\text{max\_length}}$ are then used to calculate the attention alignment map for the text encoder.
@@
$w_{\text{ceil}}$ $=(\lceil w\rceil\odot\text{speechrate})\odot\text{c\_array}$
@@ 3.1 Stimulus generation
We tested 4 different TTS styles: 1. Base: the base Matcha-TTS (0.75x speech rate), 2. Stretch: 1.2x ($0.75\times 1.6)$ speech rate applied across the entire phrase, 3. Emphasis: 1.6x stretch applied across all target words (0.75x speech rate elsewhere), and 4. Clarity: 1.6x stretch applied across the target words containing a tense vowel (0.75x speech rate elsewhere).
Previously, in [19, 22], only a single target word, always at the end of a phrase, was tested. To test the robustness of the L2 clarity TTS when the target words occurred in other parts of the phrase, we tested clarity mode in several contexts. First, single target word phrases where the target word was in the middle of the phrase were tested. We then tested double target word phrases (e.g., ``Write down dull and doll''), where the targets could be at the end of the phrase, in the middle of the phrase, or at the beginning of the phrase. Additionally, we explored the performance when the TTS had a combination of tense and lax vowels in a single phrase; we tested a tense and a lax minimal pair, a tense and a lax that are not minimal pairs, two tense and two lax vowels. We aimed to use a variety of starting and ending consonants surrounding the target vowels, and all words had a minimal pair in English. We tested all tense/lax vowel pairs: /i/ (peel) and /I/ (pill), /u/ (fool) and /U/ (full), /A/ (cot) and /2/ (cut). Lastly, we aimed not to semantically bias phrase meanings towards any word, that is, given the context outside of the target word neither word in the minimal pair made more logical or semantic sense. To do this, we asked ChatGPT-4o to provide a starting list of neutral phrases with varying lengths. These were then selected and modified to create our final list of 16 phrases and insert our target words (Table 1).
@@
Because participants would hear each of the phrases 4 times (each TTS style had the same phrases), we added 1 confusion phrase (randomly assigning a TTS style) for each of the phrases. A confusion phrase was the same phrase context using the opposite minimal pair, for example: ``She kept mentioning cot during the conversation.'' and the confusion phrase ``She kept mentioning cut during the conversation.'' Lastly, for the single target word TTS, ``clarity'' TTS was the same as ``emphasis'' for a single tense vowel target and ``base'' for a single lax vowel target. As such, we assessed the TTS in terms of the length of the target words rather than TTS styles and ensured no repeated phrases with the same treatment.
@@
iMOS 5.12 <<.001 Lax: targetS/base,fullS p=.043, p=.008; Tense: targetS/base p=.016
nMOS 110.83 <<.001 Lax: all/fullS p<<.001 targetS/base p=.006; Tense: all/fullS p<<.001
@@
pMOS 30.24 <<.001 Lax: all/fullS p<<.001 targetS/base p=.028; Tense: all/fullS p <<.001
Enc. 21.45 <<.001 Lax: all/fullS p<<.001 targetS/base p=.036; Tense: all/Stretch p<<.001
Resp. 17.90 <<.001 Lax: targetS/fullS,base p<<.001, p=0.022 base/fullS p=0.001; targetS/fullS p<<.001 base/fullS p=0.001
@@ 3.2 Experimental procedure
We conducted an experiment to assess the objective (through word error rate) and subjective (through mean opinion scores) performance of French L1 English L2 listeners using our ``clarity mode'' compared to the baseline models (i.e., ``base'', ``stretch'', ``emphasis''). In an online experiment, participants chose to receive instructions in either English or French. They then provided demographic information on the language they first learned, their most commonly used daily language, their age and gender, and their self-rated English proficiency. Each participant was randomly assigned to start in either the single- or double-word trial.
The participants were shown the phrase with the target words removed, e.g., ``Write down the word X followed by the word X on the paper,'' and could play the audio only once. They then selected which words they heard and were told (in the double-word case) that it was possible to hear the same word twice. It was never the case that they would hear the same word, however, this instruction was added to encourage participants not to base their choice for the second word on what they believe they heard in the first word. The missing word was selected from a list of 4 words: for the single-word trial, 2 words were the tense/lax vowel minimal pair (e.g., beat, bit); the third word was a minimal pair word with another vowel (e.g., bat); and the final word was dissimilar from the other three choices (e.g., shop). The dissimilar choice functioned as an attention check. In the double word case, when the two words were a minimal pair, the selection for the remaining two words was as per the single-word trial; if the two words were not a minimal pair, the 4 choices were the two words and their minimal pairs. The order of the phrases was randomized for each participant. The experiment set up can be seen in Fig. 1.
Once the participants selected which word they heard, they responded to a questionnaire containing the MOS-X2 [29] naturalness (nMOS), intelligibility (iMOS), and prosody (pMOS) scores, as well as an intelligibility question for listening effort (eMOS) (``Please rate the degree of effort you had to make to understand the message'') from MOS-X [29]. The listening effort question was added to understand if there is a difference between being able to understand the phrase and how much effort was required to understand the phrase. They then responded to two questions from L2-directed speech research [12]: ``For an English second language speaker being spoken to with this voice, how respectful is the voice?'' (Resp.: condescending - respectful), ``For an English second language speaker being spoken to with this voice, how encouraging is the voice?'' (Enc: not encouraging - encouraging). These questions were to help us understand if slowing down or adding emphasis makes the TTS sound condescending, as can be the case with L2-directed speech [11, 12]. All scores were on a 10-point Likert scale.
@@ 3.2.1 Participants
[…]
@@ 3.3.1 Single word
To confirm our previous findings, we expect improved performance in WER with lengthened tense vowels but decreased performance for the same treatment in lax vowels. We computed one-way ANOVAs (Type II) followed by a post-hoc Tukey HSD as well as word error rates (WER) as the proportion of incorrect target words for each type of target word treatment. In the single word case ``Clarity'' is the same as ``Base'' for lax vowel target words, and ``Emphasis'' for tense vowel target words. As such, we present the results in the three defined states existing for the single word case: ``base'' has the baseline duration on the target, ``target stretch'' is 1.6x the duration of the baseline on the target, and ``full stretch'' is 1.2x the duration on the entire phrase.
We observed that, through our ``clarity'' mode stretch applied to tense-vowel-containing words, we could overcome the bias towards lax vowels, i.e., the fact that participants were more likely to respond, for example, ``pill'' than ``peel,'' which we observed in our validation study. By lengthening the target tense-vowel-containing word, the WER on the baseline (``base tense'') was reduced from 60.23% to 29.48% for ``target stretch tense'' (Table 2, Fig. 2). Importantly, we also confirmed that, indeed, stretching lax vowels, as is typical in L2-directed speech, results in more errors (28.90% vs. 16.86%). Lastly, the ``stretch'' TTS resulted in a slight reduction in WER over the baseline for tense vowels, yet not as much as for only emphasizing the target word. This suggests that a difference between the target word and its context was important for determining vowel perception [19].
Surprisingly, despite the objective improvements seen in the WER, the participants perceived a ``target stretch'' significantly ($\alpha=0.05$) more intelligible for lax vowels (Tables 2 and 3, Figs. 3, 4). Moreover, ``target stretch'' also required less listening effort (although not significantly). Furthermore, although the ``full stretch'' TTS shows improvements in WER over the baseline for tense vowels, it was found to be significantly less respectful and encouraging in all cases. We also saw that the ``full stretch'' TTS was considered to be less natural and had worse prosody than the other TTS styles. This confirms our expectation that maintaining a natural (for L1 speakers) speech rate outside of difficult words can make the L2 listeners feel less patronized by the voice. Interestingly, the ``target stretch'' lax was rated as more natural and as having better prosody than ``base'' lax vowel, once again, despite have lower objective comprehension scores. These findings are particularly interesting as they suggest that L2 speakers may be overconfident in their ability to use the formant cues when the vowel is lengthened.
@@ 3.3.2 Double word
We observed the lowest WER with our proposed ``clarity'' TTS, both overall and for lax vowels, while vastly improving the performance for tense vowels over the baseline (relative 50% improvement) (Table 4, Fig. 5). Again, the ``base'' TTS showed a bias towards lax vowels (tWER: 30.00%, lWER: 18.66%), that we were able to overcome with our ``clarity'' mode. Additionally, we see that simply emphasizing all difficult words that had a tense/lax minimal pair decreases performance in lax vowels. Lastly, we observed ``stretch'' having improved performance over both ``emphasis'' and ``base'' for tense vowels. Yet, as in the ``emphasis'' condition, we saw participants struggle to understand the short, lax vowels. Although the tense vowel words between ``emphasis'' and ``clarity'' and lax vowel words between ``clarity'' and base had the same duration, ``clarity'' mode still had a lower WER when specifically comparing these words. This likely resulted from the phrases containing both a tense and a lax vowel, where the L2 participants could use duration differences between the two words to more easily differentiate the words. A more in-depth exploration of these differences remains a topic for future work.
We once again observed the fascinating result that the participants rate the ``emphasis'' TTS the highest in all categories (although in this case, the scores are not significantly higher than those for the ``clarity'' TTS) (Tables 5 and 6, and Fig. 6), despite the WER being higher for this TTS than both ``clarity'' and ``stretch'' TTS styles. We observed that ``stretch'' is less natural and has poorer prosody than all other TTS styles, despite showing objective improvements in WER over ``emphasis'' and ``base''. Lastly, we found that L2 participants rated both speaking too fast (``base'') and too slow (``stretch'') as being less respectful and less encouraging.
@@
iMOS 20.15 <<.001 all/Base p<<.001
nMOS 20.15 <<.001 all/Stretch p<<.001; Emphasis,Clarity/Base p<<.001
eMOS 24.15 <<.001 all/Base p<<.001
pMOS 77.15 <<.001 all/Stretch <<.001; Emphasis,Clarity/Base p<<.001
Enc. 35.74 <<.001 Emphasis,Clarity/Stretch,Base p<<.001
Resp. 35.87 <<.001 Emphasis,Clarity/Stretch, Base p<<.001
@@ 3.3.3 Whisper ASR
Similar to L2 speakers, we saw for the ``base'' TTS, Whisper struggled to predict the correct target words that lack context in the phrase (21.4% vs an expected 5-10% WER for this model), although the performance was slightly higher than that for the L2 speakers (Table 7, Fig 7). We did not see the same improvements in WER with the ``clarity'' TTS that we saw in the L2 participants. Instead, we saw an overall slowing down (``stretch'') of the TTS decreases the WER in the target words. Yet, we also saw that while the WER on the target words decreased with this slowing down, the proportion of errors in the target word stemming from the minimal pair substitutions was much higher for the ``base'' TTS (71.42%, all other TTS << 37%), while the overall difference in WER both on the whole phrase and the target words was within 3% for all TTS styles. This suggests that while slowing down slightly reduces the overall number of errors, the target words were being predicted as even further from the target, e.g. where ``peel'' was replaced with ``pill'' in the ``base'' TTS was replaced with ``peaked'' in the ``stretch'' TTS. Therefore, the ASR does not use the same duration mechanisms as humans when facing difficult predicting words.